OpenStudy (anonymous):

Use synthetic division to determine whether the number k is an upper or lower bound (as specified for the real zeros of the function f). k = 2; f(x) = 2x3 + 3x2 - 4x + 4; Lower bound?

OpenStudy (anonymous):

I got yes, is that correct?

OpenStudy (anonymous):

That is a different question.

OpenStudy (0_0youscareme):

oh okay

OpenStudy (anonymous):

Do you have any idea if my answer is correct?

OpenStudy (amistre64):

what does it mean to be an lower bound?

OpenStudy (anonymous):

Alternation between non negative and non positive in signs, post division.

OpenStudy (anonymous):

I think

OpenStudy (amistre64):

hmm, there is a thrm about it, can you post your materials thrm for it?

OpenStudy (amistre64):

http://www.wtamu.edu/academic/anns/mps/math/mathlab/col_algebra/col_alg_tut39_zero2.htm this seems to be what you are learning about

OpenStudy (amistre64):

2x3 + 3x2 - 4x + 4 0 4 14 20 k=2| 2 7 10 24 since all the results are the same sign ... k=2 seems to be what type of bound?

OpenStudy (amistre64):

2x3 + 3x2 - 4x + 4 0 -4 2 4 k=-2| 2 -1 -2 8 when we let k=-2, the signs are not alternating are they ...

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D2x%5E3+%2B+3x%5E2+-+4x+%2B+4 notice that it only has one real root, and it is less than 2 ... it is even less then -2 if tha tmatters

OpenStudy (anonymous):

Okay, so it is not lower bound.